### v3

# CodeWiki: Evaluating AI’s Ability to Generate Holistic Documentation for Large-Scale Codebases

Abstract
Given a large and evolving codebase, the ability to automatically generate a holistic, architecture-aware documentation that captures not only individual functions but also their cross-file, cross-module, and system-level interactions remains an open challenge. Comprehensive documentation is essential for long-term software maintenance and team collaboration, yet current automated approaches still fail to model the rich semantic dependencies and architectural structures that define real-world software systems. We present CodeWiki, a unified framework for automated repository-level documentation across seven programming languages. CodeWiki introduces three key innovations: (i) hierarchical decomposition that preserves architectural context across multiple levels of granularity, (ii) recursive multi-agent processing with dynamic task delegation for scalable generation, and (iii) multi-modal synthesis that integrates textual descriptions with visual artifacts such as architecture diagrams and data-flow representations. To enable rigorous evaluation, we introduce CodeWikiBench, a comprehensive benchmark featuring multi-dimensional rubrics and LLM-based assessment protocols. Experimental results demonstrate that CodeWiki achieves a 68.79% quality score with proprietary models, outperforming the closed-source DeepWiki baseline (64.06%) by 4.73%, with particularly strong improvements on high-level scripting languages (+10.47%). We open-source CodeWiki to foster future research and community adoption.
[…]
1 Introduction
Figure 1: CodeWiki Framework Architecture Overview. The framework operates in three main phases: (1) Repository analysis through AST/LLM parsing to construct dependency graphs and identify high-level components, followed by hierarchical decomposition into manageable modules; (2) Recursive documentation generation where specialized agents process leaf modules with dynamic delegation capabilities, creating markdown documentation while maintaining cross-module references; (3) Hierarchical assembly where parent modules are synthesized from child documentation using LLM-based synthesis, culminating in comprehensive repository overview documentation. The module tree evolves throughout the process, enabling scalable processing of repositories of arbitrary size.
In the rapidly evolving landscape of software development, maintaining comprehensive and up-to-date documentation has become increasingly challenging as codebases grow in size and complexity. Recent industry surveys reveal that documentation represents a critical bottleneck in modern development workflows: approximately 31% of developers currently rely heavily on AI for documenting code, while 25% use AI extensively for creating and maintaining documentation—making documentation tasks among the highest priorities for AI integration over the next 3-5 years [43]. Notably, while developers express high satisfaction with AI tools for searching and learning, they indicate plans to substantially increase AI adoption for documentation and testing tasks, suggesting current tools remain insufficient for these critical needs. This widespread focus on AI-driven documentation reflects its dual burden: while essential for software maintenance and collaboration [9, 58, 5], manual creation and maintenance remain labor-intensive and often neglected [30]. The emergence of Large Language Models (LLMs) [54, 17] has opened opportunities for automating documentation generation. However, current approaches still face two fundamental limitations.
Limitations of Scalability at the Repository Level
Recent advances in code understanding and natural language generation have demonstrated promising results in function-level and file-level documentation tasks [12, 20, 37, 28, 26, 1, 7, 27, 53] However, existing approaches face significant challenges when scaling to repository-level documentation.
Unlike function-level documentation that focuses on individual components, repository-level documentation task is to capture architectural patterns, cross-module interactions, data flows, and system-wide design decisions [39, 46]. Furthermore, the hierarchical nature of software systems requires documentation that can serve multiple audiences, from high-level stakeholders seeking architectural overviews to developers requiring detailed implementation guidance [31].
$\hookrightarrow$ To address these challenges, we present CodeWiki, the first open-source framework specifically designed to generate holistic, structured repository-level documentation that includes diverse content types such as system architecture diagrams, data flow visualizations, and sequence diagrams while scaling effectively for projects of arbitrary size. CodeWiki introduces several key innovations: a hierarchical decomposition strategy inspired by dynamic programming principles that breaks complex repositories into manageable modules while preserving architectural coherence, a recursive agentic system with dynamic delegation capabilities that enables adaptive processing based on module complexity, and a comprehensive documentation synthesis approach that generates multiple forms of visual and textual documentation reflecting fine-grained codebase collaboration patterns.
Limitations of the Evaluation Landscape
Despite progress, systematic benchmarks remain lacking. Function-level and file-level work relies on general-purpose metrics (BLEU, ROUGE) that fail to capture documentation quality nuances [40, 11]. Repository-level evaluation is particularly challenging as documentation admits multiple valid structures sharing underlying architectural content, highlighting the need for robust assessment beyond surface-level similarity.
$\hookrightarrow$ To fill this gap, we introduce CodeWikiBench, the first benchmark specifically designed for repository-level documentation. Instead of relying on surface-level similarity, we derive hierarchical rubrics from the official documentation maintained by project developers, enabling the evaluation of underlying repository properties. For robustness, CodeWikiBench further adopts an agentic assessment framework in which judge agents evaluate generated documentation against rubric requirements and aggregate results hierarchically, producing both quality scores and reliability measures.
Limitations of Multilingual Support
Most studies on code documentation generation have primarily focused on Python, with limited consideration of other widely used programming languages such as Java, JavaScript, C, or C++. This narrow scope restricts generalizability and overlooks the structural diversity inherent in real-world software projects, where repositories often involve multiple languages.
$\hookrightarrow$ CodeWiki supports repository-level documentation generation across 7 programming languages (Python, Java, JavaScript, TypeScript, C, C++ and C#). In addition, CodeWikiBench provides a multilingual benchmark built from multiple repositories in the same set of languages, enabling comprehensive evaluation.
Using CodeWikiBench, we demonstrate CodeWiki achieves 68.79% quality score with proprietary models, outperforming closed-source DeepWiki, while open-source alternatives achieve 64.80%, producing comprehensive documentation with architectural diagrams, usage patterns, and dependency visualizations (Appendix A).
In summary, the main contributions of this work are:
  1. 1.
    CodeWiki, the first open-source semi-agentic framework for scalable repository-level documentation with proprietary and open-source LLM backends.
  2. 2.
    CodeWikiBench, the first repository-level documentation benchmark with hierarchical rubrics and agentic assessment.
  3. 3.
    First multilingual repository documentation supporting 7 widely used languages.
  4. 4.
    Competitive performance on real-world repositories, improving over DeepWiki with both proprietary (Claude Sonnet 4) and open-source (Kimi K2 Instruct) models.
2 Related Work
2.1 Code Intelligence and LLMs
The rapid advancement of Large Language Models (LLMs) has revolutionized code intelligence tasks. Early work by von der Mosel et al. [47] demonstrated the effectiveness of Transformer-based models for software engineering, leading to specialized code models like CodeBERT [12], GraphCodeBERT [13], CodeT5 [49], CodeT5+ [50], and CodeTF [3]. These established early transformer-based approaches for code understanding and generation.
Comprehensive surveys by Niu et al. [34] and Hou et al. [19] established the importance of domain-specific pre-training, showing consistent improvements over general-purpose language models across code search, summarization, and generation tasks. However, recent studies highlight important limitations: Li et al. [22] found that models often make confident decisions without sufficient reasoning, while Zheng et al. [57] noted that LLM performance remains unstable and requires careful prompt engineering [42].
For code generation, HumanEval [6] and MBPP [2] established foundational benchmarks, later extended to multiple languages by MultiPL-E [4] and CodeScope [52]. More recently, CodeMMLU [32] has emerged as a comprehensive multi-task benchmark for assessing code understanding and reasoning capabilities of CodeLLMs.
2.2 Automated Documentation Generation
Documentation generation has emerged as a critical LLM application in software engineering. Traditional approaches focused on function-level summarization using datasets like CodeSearchNet [15] and FunCom [21]. While Sun et al. [45] found that ChatGPT can generate coherent comments, their quality often falls short of specialized models when evaluated using BLEU and ROUGE metrics.
Recent research has shifted toward repository-level documentation generation, which requires understanding complex dependencies and architectural context. Luo et al. [27] introduced RepoAgent, an LLM-powered framework that leverages static analysis and incremental processing. DocAgent [53] proposed a multi-agent system with specialized roles (Reader, Searcher, Writer, Verifier) for collaborative documentation generation. Both systems demonstrate significant improvements over baseline approaches but struggle with scaling to large repositories and maintaining cross-file consistency.
Evaluation methodologies have evolved from traditional NLP metrics like BLEU [35] and ROUGE [23] to more sophisticated approaches. Roy et al. [41] and Mastropaolo et al. [29] questioned the effectiveness of automatic metrics, while recent work has adopted LLM-as-a-judge methodologies [56, 25]. Multi-dimensional evaluation frameworks now assess completeness, helpfulness, and truthfulness, recognizing that documentation quality cannot be reduced to a single dimension.
2.3 Multi-Agent Systems for Software Engineering
Multi-agent architectures have gained prominence for tackling complex software engineering tasks requiring diverse expertise and collaborative problem-solving. AutoGen [51] introduced a framework enabling multi-agent communication with language models, allowing agents to assume different roles and collaborate on complex tasks. MetaGPT [18] demonstrated meta-programming for collaborative frameworks, where agents follow software development methodologies to produce comprehensive solutions.
Multi-agent approaches include HyperAgent [36], which presents generalist software engineering agents capable of solving coding tasks at scale, and AgileCoder [33], which introduces dynamic collaborative agents for software development based on agile methodology. In the code generation domain, MapCoder [55] employed map-reduce style collaboration to break down complex programming problems, while ChatDev [38] showed that multi-agent systems can manage the entire software development lifecycle. The integration of multi-agent systems with documentation represents a recent innovation, with DocAgent demonstrating how role-based collaboration can improve documentation quality through specialized agents for analysis, information retrieval, drafting, and verification. However, existing approaches often lack systematic dependency management and struggle with maintaining consistency across large codebases.
3 Our Approach - CodeWiki
CodeWiki is a semi-agentic framework that automatically generates comprehensive repository-level documentation by addressing context limitations through hierarchical decomposition. As illustrated in Figure 1, our approach operates through three phases: repository analysis and module decomposition, recursive documentation generation, and hierarchical assembly, inspired by dynamic programming principles.
3.1 Repository Analysis and Hierarchical Module Decomposition
We begin with static analysis to construct a dependency graph capturing structural relationships, crucial for understanding architecture and identifying decomposition boundaries.
Dependency Graph Construction
Following DocAgent [53], we employ Tree-Sitter parsers to extract ASTs, systematically identifying functions, methods, classes, structs, modules, and interdependencies: function calls, class inheritance, attribute access, and module imports. We normalize these to a unified depends_on relation for cross-language generalization, creating directed graph $G=(V,E)$ where edge $(u,v)\in E$ indicates component $u$ depends on $v$ [24].
Entry Point Identification and Decomposition
Topological sorting identifies zero-in-degree components—independent entry points where users interact (main functions, API endpoints, CLI, public interfaces). High-level components are hierarchically decomposed into manageable modules through recursive partitioning considering component interdependencies and semantic coherence. For scalability, only component IDs serve as input, resulting in a feature-oriented module tree.
3.2 Recursive Documentation Generation
CodeWiki’s core innovation is recursive agent-based processing, enabling arbitrary repository size handling while maintaining bounded complexity and architectural coherence.
Agent Architecture
Each leaf module is assigned a specialized agent equipped with: (1) complete source code access, (2) full module tree for cross-module understanding, (3) documentation workspace tools (view, create, edit operations), and (4) dependency graph traversal for contextual exploration.
The agent workflow: analyzes components to understand functionality and interfaces, explores context by traversing dependencies within and across modules, and generates comprehensive markdown documentation including descriptions, usage examples, API specifications, and architectural insights.
Dynamic Delegation
Input: Module tree $T$, dependency graph $G$
Output: Complete repository documentation
1 Initialize documentation workspace;
2 for each leaf module $m$ in $T$ do
3   agent $\leftarrow$ CreateAgent($m$, $G$, $T$);
4 while $m$ not fully processed do
5      agent.GenerateDoc($m$.name);
6    if delegation requested then
7         sub $\leftarrow$ agent.GetDelegationSpecs();
8         UpdateModuleTree($T$, $m$, sub);
9       foreach new leaf $m^{\prime}$ in sub do
10            Recursively process $m^{\prime}$;
11
12         end foreach
13        ReviseDoc($m$.name);
14
15      end if
16
17   end while
18
19 end for
20ProcessParentModules($T$);
21 return Repository documentation;
Algorithm 1 Recursive Module Documentation
Our key innovation is adaptive scalability through dynamic delegation. When module complexity exceeds single-pass capacity, agents delegate sub-modules to specialized sub-agents. Delegation criteria: code complexity metrics (cyclomatic complexity, nesting depth), semantic diversity (functionally distinct subcomponents), and context window utilization.
The recursive process follows bottom-up processing (Algorithm 1). Upon delegation, agents provide sub-module specifications, the module tree updates, and newly created leaves process recursively. This handles modules of any size while maintaining quality.
Cross-Module Reference Management
Maintaining coherence across boundaries requires sophisticated reference management. When agents encounter external components during traversal, our intelligent resolution system creates cross-references rather than duplicating content. A global registry tracks documented components and locations, enabling quick identification of documented components.
This ensures conciseness while preserving completeness, creating interconnected documentation reflecting the actual codebase structure, enabling easy navigation between related components.
3.3 Hierarchical Assembly and Documentation Synthesis
After leaf documentation, hierarchical assembly synthesizes component-level details into architectural overviews through recursive parent module processing.
Parent modules undergo LLM synthesis with carefully crafted prompting. LLMs receive: child module documentation, module tree structure, dependency information, and synthesis instructions for architectural patterns and feature interactions.
Synthesis involves multiple stages: analyzing child documentation for themes and patterns, generating architectural overviews explaining module collaboration, creating feature summaries distilling capabilities, developing usage guides for public interfaces, and producing architectural diagrams visualizing relationships and data flows.
4 The First Benchmarking System - CodeWikiBench
Repository-level documentation evaluation presents unique challenges due to complexity and hierarchical organization. Unlike code generation assessable through automated metrics, documentation quality requires nuanced evaluation of completeness, accuracy, and coherence across abstraction levels. Recent research highlights traditional metrics’ limitations for code tasks [11, 14], while LLM-as-a-Judge paradigms offer new possibilities [48, 8, 16]. We develop a comprehensive framework synthesizing repository-specific rubrics and employing agentic assessment.
4.1 Hierarchical Rubric Generation
Figure 2: Example hierarchical evaluation rubric for the RAGFlow repository. The rubric mirrors the project’s architectural structure with weighted requirements at multiple levels. Leaf nodes represent specific requirements assessed by Judge Agent, while parent scores are computed through weighted aggregation. The hierarchical organization ensures comprehensive coverage from high-level architectural components down to specific implementation details.
Following [44], we synthesize evaluation rubrics tailored for each repository, illustrated in Figure 2. We collect official documentation from selected open-source projects and hierarchically parse into structured JSON format, leveraging directory structure and markdown syntax. This captures multi-level technical documentation from architectural overviews to implementation details.
A Rubric Generator Agent processes each structure with tool access to detailed contents, generating hierarchical rubrics mirroring repository functionality and aligning with actual documentation standards. To enhance reliability, we employ multiple independent generations using different model families. Final rubrics emerge from synthesizing these perspectives, reducing single-model biases. Analysis shows generated rubrics demonstrate high consistency: 73.65% semantic reliability and 70.84% structural reliability (Appendix C).
4.2 Documentation Assessment
Our evaluation employs specialized Judge Agents operating under carefully designed protocols. Judges receive generated documentation in structured JSON format with hidden detailed contents, plus repository-specific rubrics. The systematic methodology: analyzes documentation structure and content, searches comprehensively for requirement coverage, makes binary adequacy decisions, and provides concise reasoning.
Critically, judges evaluate only leaf-level requirements, ensuring assessment grounds in specific concrete criteria rather than abstract concepts. For example, in Figure 2, agents assess "DeepDoc Visual Parser for complex layouts" rather than broader "Document Processing Engine."
For enhanced reliability, multiple Judge Agents using different model families evaluate each requirement. Final leaf scores average across assessments, reducing individual model biases.
4.3 Hierarchical Score Aggregation
Final quality scores compute through bottom-up weighted aggregation respecting rubric hierarchy while tracking reliability via standard deviation propagation. Let $R$ be the rubric root, $C(n)$ denote node $n$’s children, and $w(n)$ denote assigned weight.
Leaf Assessment and Reliability
For leaf nodes $\ell$, multiple assessments $s_{1},s_{2},\ldots,s_{m}$ aggregate as:
$S(\ell)=\bar{s}=\frac{1}{m}\sum_{i=1}^{m}s_{i}$
Reliability quantified via standard deviation:
$\sigma_{\ell}=\sqrt{\frac{1}{m-1}\sum_{i=1}^{m}(s_{i}-\bar{s})^{2}}$
Lower deviations indicate higher judge consensus and assessment reliability.
Hierarchical Propagation
For internal nodes $n$ with children $C(n)=\{c_{1},\ldots,c_{k}\}$ having scores $S(c_{i})$, weights $w(c_{i})$, and deviations $\sigma_{c_{i}}$:
$S(n)=\frac{\sum_{i=1}^{k}w(c_{i})\cdot S(c_{i})}{\sum_{i=1}^{k}w(c_{i})},\quad\sigma_{n}=\frac{\sqrt{\sum_{i=1}^{k}w(c_{i})^{2}\sigma_{c_{i}}^{2}}}{\sum_{i=1}^{k}w(c_{i})}$
This ensures uncertainty propagates appropriately, with weights determining score contribution and reliability influence.
Final Assessment
Final repository score $S(R)\in[0,1]$ with deviation $\sigma_{R}$ represents comprehensive weighted aggregation propagating assessments and reliability upward. Lower $\sigma_{R}$ indicates higher methodology reliability. This dual-metric approach provides quality assessment with confidence bounds, enabling nuanced interpretation. Weights reflect component importance based on criticality and complexity while maintaining statistical rigor in uncertainty quantification.
Repository Language LOC System Score (%) Coverage Improvement (%)
All-Hands-AI–OpenHands Python 229,909 DeepWiki 73.04 ± 2.54 54/67
CodeWiki-kimi-k2 80.20 ± 3.26 60/67 +7.16
CodeWiki-sonnet-4 82.45 ± 2.65 59/67 +9.41
sveltejs–svelte JavaScript 124,576 DeepWiki 68.51 ± 3.31 76/96
CodeWiki-kimi-k2 70.41 ± 3.39 80/96 +1.90
CodeWiki-sonnet-4 71.96 ± 3.73 80/96 +3.45
puppeteer–puppeteer TypeScript 136,302 DeepWiki 64.46 ± 3.72 60/82
CodeWiki-kimi-k2 81.33 ± 3.33 67/82 +16.87
CodeWiki-sonnet-4 83.00 ± 3.37 74/82 +18.54
Unity-Technologies–ml-agents C# 86,106 DeepWiki 74.80 ± 3.69 39/46
CodeWiki-kimi-k2 69.64 ± 5.40 40/46 -5.16
CodeWiki-sonnet-4 79.78 ± 5.02 42/46 +4.98
elastic–logstash Java 117,485 DeepWiki 54.80 ± 4.10 38/57
CodeWiki-kimi-k2 56.89 ± 2.96 37/57 +2.09
CodeWiki-sonnet-4 57.90 ± 3.43 38/57 +3.10
wazuh–wazuh C 1,446,730 DeepWiki 68.68 ± 4.74 39/46
CodeWiki-kimi-k2 58.15 ± 6.20 32/46 -10.53
CodeWiki-sonnet-4 64.17 ± 5.44 34/46 -4.51
electron–electron C++ 184,234 DeepWiki 44.10 ± 3.12 54/92
CodeWiki-kimi-k2 36.99 ± 3.98 43/92 -7.11
CodeWiki-sonnet-4 42.30 ± 3.26 48/92 -1.80
Average DeepWiki 64.06 ± 3.60
CodeWiki-kimi-k2 64.80 ± 4.07 +0.74
CodeWiki-sonnet-4 68.79 ± 3.84 +4.73
Table 1: Documentation generation performance comparison across diverse repositories. Scores represent weighted averages with standard deviations indicating assessment consensus. Coverage shows satisfied leaf criteria out of total requirements. Blue-highlighted rows show CodeWiki-sonnet-4 results, while yellow-highlighted rows show CodeWiki-kimi-k2 results. Improvement column shows relative gain over DeepWiki baseline.
5 Experiment
We conduct comprehensive experiments to evaluate CodeWiki’s effectiveness in generating repository-level documentation across diverse programming languages and project types. Our evaluation is structured around three core research questions that address the fundamental challenges of automated repository documentation generation.
5.1 Research Questions
Our experimental evaluation is guided by the following research questions:
RQ1: Documentation Quality and Coverage How does CodeWiki compare to existing state-of-the-art repository documentation systems in terms of overall documentation quality and requirement coverage?
RQ2: Cross-Language Generalization Does CodeWiki demonstrate consistent performance across diverse programming languages, particularly comparing high-level scripting languages (Python, JavaScript, TypeScript) versus systems programming languages (C, C++, C#, Java)?
RQ3: Scalability and Reliability How does CodeWiki’s hierarchical decomposition approach perform across repositories of varying sizes and complexity, and what is the reliability of our evaluation methodology?
5.2 Baseline
To the best of our knowledge, CodeWiki represents the first open-source framework specifically designed for comprehensive repository-level documentation generation. The absence of comparable open-source solutions necessitates comparison with closed-source alternatives. We select DeepWiki [10] as our primary baseline, as it shares our goal of automated repository-level documentation generation and has demonstrated effectiveness in industrial applications.
5.3 Experimental Setup
5.3.1 Repository Selection
We carefully curate a comprehensive benchmark consisting of 21 open-source repositories that represent diverse programming languages, project scales, and application domains. The selection criteria prioritize:
  1. 1.
    Language diversity to test cross-language generalization
  2. 2.
    Varying codebase sizes from 16K to 1.7M lines of code
  3. 3.
    Active maintenance with comprehensive official documentation for rubric generation
  4. 4.
    Diverse application domains including web frameworks, automation tools, machine learning platforms, and system utilities
Table 1 presents the main evaluation set of 7 repositories with detailed analysis, while Table 3 in Appendix D provides results for 14 additional repositories that further validate our approach across diverse codebases.
5.3.2 Model Configuration
We evaluate CodeWiki with two different LLM backends:
  • CodeWiki-sonnet-4: Uses Claude Sonnet 4 (proprietary model)
  • CodeWiki-kimi-k2: Uses Kimi K2 Instruct (open-source model)
For rubric generation, we employ multiple models (Claude Sonnet 4, Gemini 2.5 Pro, Kimi K2 Instruct) to reduce single-model bias. The evaluation process uses three independent judge models (Gemini 2.5 Flash, GPT OSS 120B, Kimi K2 Instruct) for robust assessment. Additional implementation details are provided in Appendix E.
5.3.3 Evaluation Protocol
For each repository, we execute the complete CodeWiki pipeline and generate comprehensive documentation including architectural overviews, component descriptions, and usage guides. The evaluation process follows our hierarchical rubric methodology:
  1. 1.
    Rubric Generation: Repository-specific rubrics derived from official documentation
  2. 2.
    Agentic Assessment: Multiple judge agents evaluate generated documentation against rubric criteria
  3. 3.
    Score Aggregation: Weighted scores computed with uncertainty quantification through standard deviation tracking
5.4 Results and Analysis
5.4.1 RQ1: Documentation Quality and Coverage
Table 1 presents comprehensive evaluation results addressing documentation quality. CodeWiki-sonnet-4 achieves an average quality score of 68.79%, representing a 4.73 percentage point improvement over DeepWiki’s 64.06%. Notably, CodeWiki-kimi-k2 demonstrates competitive performance at 64.80%, closely matching DeepWiki despite using an open-source model. This validates our framework’s effectiveness across different LLM backends.
The coverage metrics reveal that CodeWiki consistently satisfies more leaf-level requirements than the baseline. CodeWiki-sonnet-4 outperforms DeepWiki in 5 out of 7 repositories, with particularly strong improvements in TypeScript (83.00% vs. 64.46%, +18.54%), Python (82.45% vs. 73.04%, +9.41%), and JavaScript (71.96% vs. 68.51%, +3.45%). CodeWiki-kimi-k2 shows notable gains in TypeScript (+16.87%) and Python (+7.16%), demonstrating that our hierarchical approach enables strong performance even with open-source models.
Additional evaluation on 14 diverse repositories in Appendix D further validates these findings, with CodeWiki-kimi-k2 achieving an average improvement of 1.25% and substantial gains on large-scale projects including Trino (+29.53%) and Storybook (+11.42%).
Answer to RQ1: CodeWiki demonstrates superior overall documentation quality and coverage compared to existing systems. With proprietary models, it achieves 4.73% improvement; with open-source models, it matches baseline performance while offering transparency and accessibility benefits.
5.4.2 RQ2: Cross-Language Generalization
Our analysis reveals distinct performance patterns across programming language categories, as visualized in Figure 3.
Figure 3: Performance comparison by programming language categories. CodeWiki-sonnet-4 shows strong improvements in high-level scripting languages (+10.47%) and managed/enterprise languages (+4.04%), while systems programming languages present challenges for both systems.
CodeWiki-sonnet-4 demonstrates exceptional performance on high-level scripting languages, achieving 79.14% average across Python, JavaScript, and TypeScript repositories—a 10.47 percentage point improvement over DeepWiki’s 68.67%. CodeWiki-kimi-k2 achieves 77.31% on these languages, showing strong performance with open-source models (+8.64% over DeepWiki).
For managed/enterprise languages (C#, Java), CodeWiki-sonnet-4 achieves competitive performance with an average improvement of 4.04% (68.84% vs. 64.80%), while CodeWiki-kimi-k2 shows 63.27% (-1.53%). Systems programming languages present unique challenges where both frameworks exhibit similar patterns: CodeWiki-sonnet-4 achieves 53.24% compared to DeepWiki’s 56.39%, while CodeWiki-kimi-k2 obtains 47.58%. This suggests that specialized parsing strategies may benefit all approaches for handling complex constructs like pointer operations and template metaprogramming.
Answer to RQ2: CodeWiki demonstrates strong cross-language generalization with exceptional performance on high-level languages (79.14% with proprietary models, 77.31% with open-source models) and consistent improvements on managed languages, while systems programming languages present similar challenges across all frameworks and model types.
5.4.3 RQ3: Scalability and Reliability
Repository size analysis demonstrates CodeWiki’s scalability across 86K-1.4M LOC. For high-level languages, CodeWiki maintains strong performance across scales: Unity ML-Agents (86K LOC) achieves 79.78%, while OpenHands (230K LOC) reaches 82.45%, demonstrating that hierarchical decomposition successfully addresses context limitations. Both CodeWiki and DeepWiki exhibit similar behavioral patterns across repository sizes within language families, with performance variations primarily attributable to language-specific parsing complexity rather than scale. Dynamic delegation automatically adapts to complexity levels, maintaining bounded processing requirements.
Reliability analysis through standard deviations (3.84% CodeWiki vs. 3.60% DeepWiki) validates our multi-model assessment methodology. Lower deviations for high-level languages (Python: 2.65%, JavaScript: 3.73%) versus systems languages (C: 5.44%, C++: 3.26%) confirm evaluation consistency correlates with language characteristics rather than repository scale.
Answer to RQ3: CodeWiki’s hierarchical decomposition enables consistent scalability across repository sizes from 86K to 1.4M LOC, with performance variations primarily attributable to language-specific characteristics rather than scale limitations. Both frameworks exhibit similar patterns across different repository sizes within language families.
6 Conclusion
We presented CodeWiki, an open-source semi-agentic framework addressing automated repository-level documentation through hierarchical decomposition and recursive processing. CodeWiki achieves 68.79% quality with proprietary models (Claude Sonnet 4) and 64.80% with open-source alternatives (Kimi K2 Instruct), both outperforming or matching DeepWiki’s 64.06%, demonstrating strong performance regardless of underlying model choice.
Results show performance variations correlate primarily with language characteristics rather than repository size or model type. High-level languages show consistent improvements across scales and models, while systems languages face challenges regardless of configuration.
Future directions include: developing specialized parsing modules for improved systems language support, extending to multi-version documentation tracking codebase evolution, and leveraging comprehensive documentation for downstream tasks like code migration. By releasing CodeWiki as open source with support for both proprietary and open-source LLMs, we accelerate research and enable the community to build upon our hierarchical processing approach with accessible technology choices.
Limitations
While CodeWiki demonstrates strong performance across multiple languages and repository scales, two main limitations warrant discussion. First, our framework exhibits reduced effectiveness on systems programming languages (C, C++) compared to high-level languages, achieving 53.24% versus 79.14% average scores. This gap stems from the inherent complexity of low-level constructs such as pointer operations, manual memory management, and template metaprogramming, which challenge both AST parsing and semantic understanding.
Second, our evaluation methodology, while more comprehensive than traditional metrics, relies on LLM-based assessment which may introduce model-specific biases despite our multi-model consensus approach. The average standard deviation of 3.84% indicates reasonable consistency, yet edge cases may benefit from human expert validation.
Appendices
Appendix A Example Generated Documentation
Figure 4: Example repository-level documentation generated by CodeWiki for the All-Hands-AI–OpenHands repository. The documentation includes a comprehensive overview, architectural diagrams showing the modular event-driven architecture, and hierarchical navigation of core components including the agent system, implementations, and event processing modules.
Figure 4 shows an example of repository-level documentation generated by CodeWiki for the All-Hands-AI–OpenHands repository. The documentation provides a comprehensive overview including the repository’s purpose, end-to-end architecture with visual diagrams, and hierarchical navigation of core components. This demonstrates CodeWiki’s ability to synthesize both textual explanations and visual artifacts that capture architectural patterns and module relationships, enabling developers to quickly understand the system’s structure and design decisions.
Appendix B Related Work
B.1 Code Intelligence and LLMs
The rapid advancement of Large Language Models (LLMs) has revolutionized code intelligence tasks. Early work by von der Mosel et al. [47] demonstrated the effectiveness of Transformer-based models for software engineering, leading to specialized code models like CodeBERT [12], GraphCodeBERT [13], CodeT5 [49], CodeT5+ [50], and CodeTF [3]. These established early transformer-based approaches for code understanding and generation.
Comprehensive surveys by Niu et al. [34] and Hou et al. [19] established the importance of domain-specific pre-training, showing consistent improvements over general-purpose language models across code search, summarization, and generation tasks. However, recent studies highlight important limitations: Li et al. [22] found that models often make confident decisions without sufficient reasoning, while Zheng et al. [57] noted that LLM performance remains unstable and requires careful prompt engineering [42].
For code generation, HumanEval [6] and MBPP [2] established foundational benchmarks, later extended to multiple languages by MultiPL-E [4] and CodeScope [52]. More recently, CodeMMLU [32] has emerged as a comprehensive multi-task benchmark for assessing code understanding and reasoning capabilities of CodeLLMs.
B.2 Automated Documentation Generation
Documentation generation has emerged as a critical LLM application in software engineering. Traditional approaches focused on function-level summarization using datasets like CodeSearchNet [15] and FunCom [21]. While Sun et al. [45] found that ChatGPT can generate coherent comments, their quality often falls short of specialized models when evaluated using BLEU and ROUGE metrics.
Recent research has shifted toward repository-level documentation generation, which requires understanding complex dependencies and architectural context. Luo et al. [27] introduced RepoAgent, an LLM-powered framework that leverages static analysis and incremental processing. DocAgent [53] proposed a multi-agent system with specialized roles (Reader, Searcher, Writer, Verifier) for collaborative documentation generation. Both systems demonstrate significant improvements over baseline approaches but struggle with scaling to large repositories and maintaining cross-file consistency.
Evaluation methodologies have evolved from traditional NLP metrics like BLEU [35] and ROUGE [23] to more sophisticated approaches. Roy et al. [41] and Mastropaolo et al. [29] questioned the effectiveness of automatic metrics, while recent work has adopted LLM-as-a-judge methodologies [56, 25]. Multi-dimensional evaluation frameworks now assess completeness, helpfulness, and truthfulness, recognizing that documentation quality cannot be reduced to a single dimension.
B.3 Multi-Agent Systems for Software Engineering
Multi-agent architectures have gained prominence for tackling complex software engineering tasks requiring diverse expertise and collaborative problem-solving. AutoGen [51] introduced a framework enabling multi-agent communication with language models, allowing agents to assume different roles and collaborate on complex tasks. MetaGPT [18] demonstrated meta-programming for collaborative frameworks, where agents follow software development methodologies to produce comprehensive solutions.
Multi-agent approaches include HyperAgent [36], which presents generalist software engineering agents capable of solving coding tasks at scale, and AgileCoder [33], which introduces dynamic collaborative agents for software development based on agile methodology. In the code generation domain, MapCoder [55] employed map-reduce style collaboration to break down complex programming problems, while ChatDev [38] showed that multi-agent systems can manage the entire software development lifecycle. The integration of multi-agent systems with documentation represents a recent innovation, with DocAgent demonstrating how role-based collaboration can improve documentation quality through specialized agents for analysis, information retrieval, drafting, and verification. However, existing approaches often lack systematic dependency management and struggle with maintaining consistency across large codebases.
Appendix C Rubric Reliability Assessment
To ensure the quality and trustworthiness of automatically generated evaluation rubrics, we developed a comprehensive reliability assessment framework that quantifies the consistency and structural integrity of rubrics produced by different language models. This assessment methodology provides empirical measures of rubric reliability across multiple dimensions.
Given a set of $n$ rubrics $\{R_{1},R_{2},\ldots,R_{n}\}$ produced by different models for the same documentation corpus, we compute pairwise consistency scores across both semantic and structural dimensions.
C.1 Semantic Consistency Analysis
We evaluate semantic consistency by extracting all requirement texts from each rubric and compute semantic similarity using distributed representations. Let $T_{i}=\{t_{1}^{(i)},t_{2}^{(i)},\ldots,t_{m_{i}}^{(i)}\}$ denote the set of requirement texts extracted from rubric $R_{i}$, where $m_{i}$ represents the total number of requirements in $R_{i}$.
For each requirement text $t_{k}^{(i)}\in T_{i}$, we identify its best semantic match in $T_{j}$ using cosine similarity between embeddings:
$\text{sim}_{\text{best}}(t_{k}^{(i)},T_{j})=\max_{t_{\ell}^{(j)}\in T_{j}}\cos(\mathbf{e}(t_{k}^{(i)}),\mathbf{e}(t_{\ell}^{(j)}))$ (1)
where $\mathbf{e}(\cdot)$ represents the embedding function and $\cos(\cdot,\cdot)$ denotes cosine similarity. The bidirectional semantic similarity between rubrics is then:
$\text{Sim}_{\text{semantic}}(R_{i},R_{j})$ $=\frac{1}{m_{i}+m_{j}}\Bigg(\sum_{k=1}^{m_{i}}\text{sim}_{\text{best}}(t_{k}^{(i)},T_{j})$
$\quad+\sum_{\ell=1}^{m_{j}}\text{sim}_{\text{best}}(t_{\ell}^{(j)},T_{i})\Bigg)$ (2)
This formulation ensures that both the coverage of requirements from $R_{i}$ in $R_{j}$ and vice versa are considered, providing a balanced measure of semantic overlap.
C.2 Structural Consistency Analysis
Structural consistency evaluates the architectural similarity between rubrics, focusing on hierarchical organization rather than content semantics. We define a comprehensive set of structural features for each rubric:
  • Maximum depth ($d$): The deepest level in the rubric hierarchy
  • Total items ($N$): The total number of rubric items across all levels
  • Weight distribution ($W$): The frequency distribution of importance weights
The structural similarity between rubrics $R_{i}$ and $R_{j}$ combines multiple normalized metrics:
$\text{Sim}_{\text{depth}}(R_{i},R_{j})=1-\frac{|d_{i}-d_{j}|}{\max(d_{i},d_{j},1)}$ (3)
$\text{Sim}_{\text{items}}(R_{i},R_{j})=1-\frac{|N_{i}-N_{j}|}{\max(N_{i},N_{j},1)}$ (4)
For weight distribution similarity, we compute the overlap between normalized probability distributions:
$\text{Sim}_{\text{weights}}(R_{i},R_{j})=\sum_{w\in W_{i}\cup W_{j}}\min(P_{i}(w),P_{j}(w))$ (5)
where $P_{i}(w)$ represents the normalized frequency of weight $w$ in rubric $R_{i}$.
The overall structural similarity is computed as:
$\text{Sim}_{\text{structural}}(R_{i},R_{j})$ $=\frac{1}{3}\Big(\text{Sim}_{\text{depth}}(R_{i},R_{j})$
$\quad+\text{Sim}_{\text{items}}(R_{i},R_{j})$ (6)
$\quad+\text{Sim}_{\text{weights}}(R_{i},R_{j})\Big)$ (7)
C.3 Overall Reliability Score
The overall reliability score aggregates consistency measures across all model pairs. For $n$ models, we compute $\binom{n}{2}$ pairwise comparisons and derive summary statistics:
$\text{Score}^{semantic}_{\text{reliability}}=\frac{1}{\binom{n}{2}}\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}\text{Sim}_{\text{semantic}}(R_{i},R_{j})$ (8)
$\text{Score}^{structural}_{\text{reliability}}=\frac{1}{\binom{n}{2}}\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}\text{Sim}_{\text{structural}}(R_{i},R_{j})$ (9)
Table 2 presents the reliability assessment results across our benchmark repositories, demonstrating consistent rubric generation with average semantic reliability of 73.65% and structural reliability of 70.84%.
Table 2: Rubric Reliability Assessment: Semantic and Structural Consistency Scores
Repository Semantic (%) Structural (%)
All-Hands-AI–OpenHands 73.64 68.82
sveltejs–svelte 75.56 68.88
puppeteer–puppeteer 72.97 76.36
Unity-Technologies–ml-agents 75.95 71.22
elastic–logstash 73.13 68.22
wazuh–wazuh 72.20 67.62
electron–electron 72.12 74.79
Average 73.65 70.84
Appendix D Complementary Experimental Results
To further validate CodeWiki’s effectiveness across diverse programming languages and repository types, we conducted additional experiments on 14 repositories spanning 7 programming languages. Table 3 presents comprehensive results using CodeWiki-kimi-k2 (documentation generated using Kimi K2 Instruct) compared against the DeepWiki baseline.
Repository Lang. LOC System Score (%) Cov. Imp. (%)
chartjs–Chart.js JavaScript 90,925 DeepWiki 88.71 ± 3.29 57/61
CodeWiki-kimi-k2 84.01 ± 3.55 54/61 -4.70
marktext–marktext JavaScript 61,574 DeepWiki 81.64 ± 4.67 50/57
CodeWiki-kimi-k2 82.49 ± 2.51 51/57 +0.85
storybookjs–storybook TypeScript 364,606 DeepWiki 72.07 ± 3.59 64/78
CodeWiki-kimi-k2 83.49 ± 2.31 69/78 +11.42
mermaid-js–mermaid TypeScript 157,160 DeepWiki 73.81 ± 4.05 59/67
CodeWiki-kimi-k2 71.52 ± 5.36 57/67 -2.29
qmk–qmk_firmware C 1,099,824 DeepWiki 81.23 ± 2.62 94/110
CodeWiki-kimi-k2 72.01 ± 2.63 88/110 -9.22
sumatrapdfreader–sumatrapdf C 1,699,044 DeepWiki 55.81 ± 6.29 31/40
CodeWiki-kimi-k2 58.92 ± 5.18 31/40 +3.11
x64dbg–x64dbg C++ 207,671 DeepWiki 64.42 ± 2.07 62/107
CodeWiki-kimi-k2 58.01 ± 3.58 56/107 -6.41
nlohmann–json C++ 119,542 DeepWiki 66.06 ± 3.08 33/57
CodeWiki-kimi-k2 61.28 ± 2.35 30/57 -4.78
FluentValidation–FluentValidation C# 16,452 DeepWiki 76.48 ± 2.33 41/48
CodeWiki-kimi-k2 70.03 ± 2.90 38/48 -6.45
git-ecosystem–git-credential-manager C# 40,812 DeepWiki 87.03 ± 2.58 66/70
CodeWiki-kimi-k2 82.78 ± 0.90 61/70 -4.25
material-components–material-components-android Java 244,092 DeepWiki 70.99 ± 2.92 79/96
CodeWiki-kimi-k2 79.16 ± 4.32 89/96 +8.17
trinodb–trino Java 1,535,202 DeepWiki 55.27 ± 3.76 48/68
CodeWiki-kimi-k2 84.80 ± 2.84 62/68 +29.53
RasaHQ–rasa Python 147,711 DeepWiki 80.94 ± 3.15 64/70
CodeWiki-kimi-k2 81.10 ± 2.69 64/70 +0.16
microsoft–graphrag Python 30,390 DeepWiki 76.34 ± 3.28 48/56
CodeWiki-kimi-k2 79.29 ± 2.96 50/56 +2.95
Average DeepWiki 73.62 ± 3.41
CodeWiki-kimi-k2 74.87 ± 3.15 +1.25
Table 3: Complementary documentation generation performance evaluation across 14 additional repositories. Results demonstrate CodeWiki’s consistent effectiveness across diverse codebases, with notable improvements in Java repositories (trinodb–trino: +29.53%, material-components-android: +8.17%) and TypeScript (storybookjs–storybook: +11.42%).
These complementary results reinforce our findings from the main experiments, showing an average improvement of 1.25% across 14 diverse repositories. Notably, CodeWiki achieves substantial gains in several large-scale projects, including a remarkable 29.53% improvement on the Trino database engine (1.5M LOC) and 11.42% on the Storybook UI development platform (365K LOC). The results further validate that CodeWiki’s hierarchical approach generalizes well across different repository characteristics and programming paradigms.
Appendix E Implementation Details
E.1 Model Configurations
Documentation Generation Agent
We evaluate CodeWiki with two model configurations:
  • CodeWiki-sonnet-4: Claude Sonnet 4 (claude-sonnet-4-20250514)
    • Temperature: 0.0
    • Max tokens: 32768
    • Context window: 200K tokens
  • CodeWiki-kimi-k2: Kimi K2 Instruct (kimi-k2-instruct-0905)
    • Temperature: 0.0
    • Max tokens: 32768
    • Context window: 200K tokens
Rubric Generation Agents
  • Model 1: Claude Sonnet 4 (Temperature: 0.0)
  • Model 2: Gemini 2.5 Pro (Temperature: 0.0)
  • Model 3: Kimi K2 Instruct (Temperature: 0.0)
  • Max tokens per generation: 32768
Judge Agents
  • Judge 1: Gemini 2.5 Flash (Temperature: 0.0)
  • Judge 2: GPT OSS 120B (Temperature: 0.0)
  • Judge 3: Kimi K2 Instruct (Temperature: 0.0)
  • Binary scoring: 0 (not satisfied) or 1 (satisfied)
E.2 Processing Parameters
  • Module decomposition threshold: 32768 tokens per leaf module
  • Maximum delegation depth: 3 levels