### v5

# CodeWiki: Evaluating AI’s Ability to Generate Holistic Documentation for Large-Scale Codebases

Abstract
Comprehensive software documentation is crucial yet costly to produce. Despite recent advances in large language models (LLMs), generating holistic, architecture-aware documentation at the repository level remains challenging due to complex and evolving codebases that exceed LLM context limits. Existing automated methods struggle to capture rich semantic dependencies and architectural structure. We present CodeWiki, a unified framework for automated repository-level documentation across seven mainstream programming languages. CodeWiki combines top-down hierarchical decomposition with a divide-and-conquer agent system to preserve architectural context and scale documentation generation, and a bottom-up synthesis that integrates textual descriptions with visual artifacts such as architecture and data-flow diagrams. We also introduce CodeWikiBench, a benchmark with hierarchical rubrics and LLM-based evaluation protocols. Experiments show that CodeWiki achieves a 68.79% quality score with proprietary models, outperforming the closed-source DeepWiki baseline by 4.73%, with especially strong gains on scripting languages. CodeWiki is released as open source to support future research.
1 Introduction
Figure 1: CodeWiki Framework Architecture Overview. The framework operates in three main phases: (1) Repository analysis through AST/LLM parsing to construct dependency graphs and identify high-level components, followed by hierarchical decomposition into manageable modules; (2) Recursive documentation generation where specialized agents process leaf modules with dynamic delegation capabilities, creating markdown documentation while maintaining cross-module references; (3) Hierarchical assembly where parent modules are synthesized from child documentation using LLM-based synthesis, culminating in comprehensive repository overview documentation. The module tree evolves throughout the process, enabling scalable processing of repositories of arbitrary size.
In the rapidly evolving landscape of software development, maintaining comprehensive and up-to-date documentation has become increasingly challenging as codebases grow in size and complexity. Studies indicate that developers spend approximately 58% of their working time on program comprehension activities (Xia et al., 2018), highlighting the critical importance of accessible and accurate documentation. Despite documentation being recognized as essential for software maintenance and collaboration (de Souza et al., 2005; Zhi et al., 2015; Chen and Huang, 2009), the manual creation and maintenance of repository-level documentation remain a labor-intensive and often neglected aspect of software engineering practice (McBurney et al., 2018). The emergence of Large Language Models (LLMs)  (Vaswani et al., 2017; Zhang et al., 2023; He et al., 2025b) has opened opportunities for automating documentation generation tasks. However, current approaches still face two fundamental limitations.
Limitations of Scalability at the Repository Level.
Recent advances in code understanding and natural language generation have demonstrated promising results in function and file documentation tasks (Feng et al., 2020; Khan and Uddin, 2023; Poudel et al., 2024; Makharev and Ivanov, 2025; Lomshakov et al., 2024; Ahmed and Devanbu, 2023; Choi et al., 2023; Luo et al., 2024; Yang et al., 2025). However, existing approaches face significant challenges when scaling to repository documentation. Unlike function documentation that focuses on individual components, repository documentation must capture architectural patterns, interactions between modules, data flows, and design decisions that span the entire system (Rai et al., 2022; Treude et al., 2020). Furthermore, the complex hierarchical nature of software systems requires documentation that can serve multiple purposes, such as for stakeholders seeking architectural overviews or developers requiring detailed implementation guidance (Nassif and Robillard, 2025).
$\hookrightarrow$ To address these challenges, we present CodeWiki, an open-source framework designed to generate holistic repository documentation that includes diverse content types such as system architecture diagrams, data flow visualizations, and sequence diagrams while scaling effectively for projects of arbitrary size. While prior work such as RepoAgent (Luo et al., 2024) aggregates function and class documentation into repository structures, and open-source alternatives like Open-DeepWiki apply LLMs to entire repositories, CodeWiki differs through its hierarchical synthesis approach where parent modules are synthesized from child documentation, combined with integrated visual artifacts capturing architectural patterns. CodeWiki introduces three key innovations: (1) a hierarchical decomposition strategy inspired by dynamic programming principles that breaks complex repositories into manageable modules while preserving architectural coherence; (2) a scalable divide-and-conquer agent system that dynamically determines the relevant context for each module, recursively partitions it into sub-modules handled by dedicated sub-agents; and (3) a bottom-up synthesis process that integrates textual descriptions with visual artifacts such as architecture diagrams and data-flow representations to generate high-level documentation.
Limitations of the Evaluation Landscape.
Despite progress in automated documentation, there is still a lack of systematic and standardized benchmarks. Existing work at the function-level and file-level has mostly been evaluated using general-purpose text generation metrics (e.g., BLEU, ROUGE), which fail to capture the nuanced quality of software documentation (Roy et al., 2021; Evtikhiev et al., 2023). For repository-level documentation, the large scale and complexity of software systems make evaluation particularly challenging, as the quality of documentation cannot be adequately judged by surface-level text-similarity metrics alone. A single repository can be described through documentation with different structures and presentations, yet all share the same underlying architectural content. This highlights the lack of a robust evaluation system that can fairly assess documentation quality regardless of how it is presented.
$\hookrightarrow$ To fill this gap, we introduce CodeWikiBench, the first benchmark specifically designed for repository-level documentation. Instead of relying on surface-level similarity, we automatically derive hierarchical evaluation rubrics from the ground-truth official documentation maintained by project developers, enabling the automated evaluation of machine-generated documentation using the derived rubrics. This approach is inspired by recent advances in hierarchical rubric-based evaluation (Starace et al., 2025), which demonstrates the effectiveness of decomposing complex evaluation tasks into smaller, well-defined sub-tasks with clear grading criteria. For robustness, CodeWikiBench adopts an agentic assessment framework in which judge agents evaluate generated documentation against rubric requirements and aggregate results hierarchically, producing both quality scores and reliability measures.
Limitations of Multilingual Support.
Most studies on code documentation generation have primarily focused on Python, with limited consideration of other widely used programming languages such as Java, JavaScript, C, or C++. This narrow scope restricts generalizability and overlooks the structural diversity inherent in real-world software projects, where repositories often involve multiple languages.
$\hookrightarrow$ CodeWiki supports repository-level documentation generation across 7 programming languages (Python, Java, JavaScript, TypeScript, C, C++ and C#). In addition, CodeWikiBench provides a multilingual benchmark built from multiple repositories in the same set of languages, enabling comprehensive evaluation.
Using CodeWikiBench, we evaluate the effectiveness of CodeWiki against multiple repository-level documentation systems, including open-source baselines (OpenDeepWiki [^1] https://github.com/AIDotNet/OpenDeepWiki, deepwiki-open [^2] https://github.com/AsyncFuncAI/deepwiki-open) and the closed-source DeepWiki. The results show that CodeWiki achieves an overall quality score of 68.79%, outperforming all baselines with improvements of up to 18.54% over the strongest baseline, DeepWiki, on individual repositories. CodeWiki produces comprehensive documentation encompassing architectural diagrams, usage patterns, and cross-module dependency visualizations (see Appendix A for an example).
In summary, our main contributions are:
  1. 1.
    CodeWiki, an open-source framework for scalable repository documentation generation combining top-down and bottom-up analyses.
  2. 2.
    CodeWikiBench, a benchmark for repository documentation featuring hierarchical rubrics and an agentic assessment framework.
  3. 3.
    Support for documentation generation across seven programming languages (Python, Java, JavaScript, TypeScript, C, C++, and C#).
  4. 4.
    Experimental validation demonstrating improvements over all baselines, with gains of up to 18.54% on individual repositories and 4.73% on average compared to the closed-source DeepWiki system.
A detailed discussion of related work and positioning relative to prior approaches is provided in Appendix B.
2 Our Approach - CodeWiki
CodeWiki is a semi-agentic framework that automatically generates comprehensive repository-level documentation by addressing context limitations through hierarchical decomposition. As illustrated in Figure 1, our approach operates through three phases: repository analysis and module decomposition, recursive documentation generation, and hierarchical assembly, inspired by dynamic programming principles.
2.1 Repository Analysis and Hierarchical Module Decomposition
We begin with static analysis to construct a dependency graph capturing structural relationships, crucial for understanding architecture and identifying decomposition boundaries.
Dependency Graph Construction
Following DocAgent (Yang et al., 2025), we employ Tree-Sitter parsers to extract Abstract Syntax Trees (ASTs). We systematically identify functions, methods, classes, structs, modules, and interdependencies including: function calls (runtime dependencies), class inheritance (structural hierarchies), attribute access (data dependencies), and module imports (compilation dependencies). We normalize these to a unified depends_on relation for cross-language generalization.
This creates a directed graph $G=(V,E)$ where vertices $V$ represent components and edge $(u,v)\in E$ indicates component $u$ depends on $v$. Understanding $v$ is prerequisite to comprehending $u$, fundamental to our hierarchical processing (Liu et al., 2023).
Entry Point Identification and Decomposition
Topological sorting identifies zero-in-degree components, independent entry points where users interact with the project (main functions, API endpoints, CLI, public interfaces). Components from these files represent high-level features and documentation starting points.
High-level components are hierarchically decomposed into manageable modules through recursive partitioning. For scalability, only component IDs (relative paths) serve as input. Decomposition considers: component interdependencies (grouping related functionality) and semantic coherence (preserving logical boundaries). The result is a feature-oriented module tree where each node contains component lists and sub-modules.
2.2 Recursive Documentation Generation
CodeWiki’s core innovation is recursive agent processing, enabling arbitrary repository size handling while maintaining bounded complexity and architectural coherence.
Agent Architecture
Each leaf module is assigned a specialized agent equipped with: (1) complete source code access, (2) full module tree for cross-module understanding, (3) documentation workspace tools, and (4) dependency graph traversal for contextual exploration.
The agent workflow: analyzes components to understand functionality and interfaces, explores context by traversing dependencies, and generates comprehensive markdown documentation including descriptions, usage examples, API specifications, and architectural insights.
Dynamic Delegation
Our key innovation is adaptive scalability through dynamic delegation. When module complexity exceeds single-pass capacity, agents delegate submodules to specialized subagents. Delegation criteria include code complexity metrics, semantic diversity, and context window utilization.
The recursive process follows bottom-up processing (see Algorithm 1 in Appendix C). Upon delegation, agents provide submodule specifications, the module tree updates, and newly created leaves process recursively. This enables handling modules of any size while maintaining quality.
Cross-Module Reference Management
Maintaining coherence across boundaries requires sophisticated reference management. When agents encounter external components, our resolution system creates cross-references rather than duplicating content. A global registry tracks documented components, enabling quick identification and navigation between related components.
2.3 Hierarchical Assembly and Documentation Synthesis
After leaf documentation, hierarchical assembly synthesizes component-level details into architectural overviews through recursive parent module processing.
Parent modules undergo LLM synthesis with carefully crafted prompting. LLMs receive: child module documentation, module tree structure, dependency information, and synthesis instructions for architectural patterns and feature interactions.
Synthesis involves multiple stages: analyzing child documentation for themes and patterns, generating architectural overviews explaining module collaboration, creating feature summaries distilling capabilities, developing usage guides for public interfaces, and producing architectural diagrams visualizing relationships and data flows.
3 The First Benchmarking System - CodeWikiBench
Figure 2: Example hierarchical evaluation rubric for the RAGFlow repository. The rubric mirrors the project’s architectural structure with weighted requirements at multiple levels. Leaf nodes represent specific requirements assessed by Judge Agent, while parent scores are computed through weighted aggregation. The hierarchical organization ensures comprehensive coverage from high-level architectural components down to specific implementation details, with weights reflecting the relative importance of each component for understanding the repository
Repository-level documentation evaluation presents unique challenges due to complexity and hierarchical organization. Unlike code generation assessable through automated metrics, documentation quality requires nuanced evaluation of completeness, accuracy, and coherence across abstraction levels. Recent research highlights traditional metrics’ limitations for code tasks (Evtikhiev et al., 2023; Haldar and Hockenmaier, 2024), while LLM-as-a-Judge paradigms offer new possibilities (Wang et al., 2025; Crupi et al., 2025; He et al., 2025a). We develop a comprehensive framework synthesizing repository-specific rubrics and employing agentic assessment.
3.1 Hierarchical Rubric Generation
Following (Starace et al., 2025), we synthesize evaluation rubrics tailored for each repository, illustrated in Figure 2. We collect official documentation from selected open-source projects and hierarchically parse into structured JSON format, leveraging directory structure and markdown syntax.
A Rubric Generator Agent processes each structure with tool access to detailed contents, generating hierarchical rubrics mirroring repository functionality. To enhance reliability, we employ multiple independent generations using different model families. Final rubrics emerge from synthesizing these perspectives, reducing single-model biases. Analysis shows generated rubrics demonstrate high consistency: 73.65% semantic reliability and 70.84% structural reliability (Appendix D).
3.2 Documentation Assessment
Our evaluation employs specialized Judge Agents operating under carefully designed protocols. Judges receive generated documentation in structured JSON format with hidden detailed contents, plus repository-specific rubrics. The systematic methodology: analyzes documentation structure and content, searches comprehensively for requirement coverage, makes binary adequacy decisions, and provides concise reasoning.
Critically, judges evaluate only leaf-level requirements, ensuring assessment grounds in specific concrete criteria rather than abstract concepts. For example, in Figure 2, agents assess "DeepDoc Visual Parser for complex layouts" rather than broader "Document Processing Engine."
For enhanced reliability, multiple Judge Agents using different model families evaluate each requirement. Final leaf scores average across assessments, reducing individual model biases.
3.3 Hierarchical Score Aggregation
Final quality scores compute through bottom-up weighted aggregation respecting rubric hierarchy while tracking reliability via standard deviation propagation. Let $R$ be the rubric root, $C(n)$ denote node $n$’s children, and $w(n)$ denote assigned weight.
Leaf Assessment and Reliability
For leaf nodes $\ell$, multiple assessments $s_{1},s_{2},\ldots,s_{m}$ aggregate as:
$S(\ell)=\bar{s}=\frac{1}{m}\sum_{i=1}^{m}s_{i}$
Reliability quantified via standard deviation:
$\sigma_{\ell}=\sqrt{\frac{1}{m-1}\sum_{i=1}^{m}(s_{i}-\bar{s})^{2}}$
Lower deviations indicate higher judge consensus and assessment reliability.
Hierarchical Propagation
For internal nodes $n$ with children $C(n)=\{c_{1},\ldots,c_{k}\}$ having scores $S(c_{i})$, weights $w(c_{i})$, and deviations $\sigma_{c_{i}}$:
$S(n)=\frac{\sum_{i=1}^{k}w(c_{i})\cdot S(c_{i})}{\sum_{i=1}^{k}w(c_{i})}$
$\sigma_{n}=\frac{\sqrt{\sum_{i=1}^{k}w(c_{i})^{2}\sigma_{c_{i}}^{2}}}{\sum_{i=1}^{k}w(c_{i})}$
This ensures uncertainty propagates appropriately, with weights determining score contribution and reliability influence.
Final Assessment
Final repository score $S(R)\in[0,1]$ with deviation $\sigma_{R}$ represents weighted aggregation propagating assessments and reliability upward. Lower $\sigma_{R}$ indicates higher methodology reliability. This dual-metric approach provides quality assessment with confidence bounds, enabling nuanced interpretation. Weights reflect component importance based on criticality and complexity while maintaining statistical rigor in uncertainty quantification.
Repository Language LOC System Score (%) Coverage Improvement (%)
All-Hands-AI–OpenHands Python 229,909 OpenDeepWiki 58.12 ± 3.21 42/67
deepwiki-open 61.35 ± 2.98 45/67
DeepWiki 73.04 ± 2.54 54/67
CodeWiki 82.45 ± 2.65 59/67 +9.41
sveltejs–svelte JavaScript 124,576 OpenDeepWiki 55.23 ± 3.85 61/96
deepwiki-open 57.89 ± 3.62 64/96
DeepWiki 68.51 ± 3.31 76/96
CodeWiki 71.96 ± 3.73 80/96 +3.45
puppeteer–puppeteer TypeScript 136,302 OpenDeepWiki 51.82 ± 4.15 48/82
deepwiki-open 54.67 ± 3.94 51/82
DeepWiki 64.46 ± 3.72 60/82
CodeWiki 83.00 ± 3.37 74/82 +18.54
Unity-Technologies–ml-agents C# 86,106 OpenDeepWiki 62.45 ± 4.28 32/46
deepwiki-open 65.12 ± 4.05 34/46
DeepWiki 74.80 ± 3.69 39/46
CodeWiki 79.78 ± 5.02 42/46 +4.98
elastic–logstash Java 117,485 OpenDeepWiki 41.25 ± 4.52 28/57
deepwiki-open 44.18 ± 4.31 31/57
DeepWiki 54.80 ± 4.10 38/57
CodeWiki 57.90 ± 3.43 38/57 +3.10
wazuh–wazuh C 1,446,730 OpenDeepWiki 32.56 ± 5.82 18/46
deepwiki-open 35.89 ± 5.45 21/46
DeepWiki 68.68 ± 4.74 39/46
CodeWiki 64.17 ± 5.44 34/46 -4.51
electron–electron C++ 184,234 OpenDeepWiki 28.45 ± 3.95 35/92
deepwiki-open 31.22 ± 3.78 38/92
DeepWiki 44.10 ± 3.12 54/92
CodeWiki 42.30 ± 3.26 48/92 -1.80
Average OpenDeepWiki 47.13 ± 4.25
deepwiki-open 50.05 ± 4.02
DeepWiki 64.06 ± 3.60
CodeWiki 68.79 ± 3.84 +4.73
Table 1: Documentation generation performance comparison across repositories. Scores represent weighted averages with standard deviations indicating assessment consensus. Coverage shows satisfied leaf criteria out of total requirements. OpenDeepWiki, deepwiki-open are open-source alternatives; DeepWiki is the closed-source baseline.
4 Experiment
We conduct comprehensive experiments to assess CodeWiki’s effectiveness in generating repository-level documentation across diverse programming languages and project types. Our evaluation is structured around 3 core research questions that address the fundamental challenges of automated repository documentation generation.
4.1 Research Questions
Our experimental evaluation is guided by the following research questions:
RQ1: Documentation Quality and Coverage How does CodeWiki compare to existing state-of-the-art repository documentation systems in terms of overall documentation quality and requirement coverage?
RQ2: Language Generalization Does CodeWiki demonstrate consistent performance across diverse programming languages, particularly comparing high-level scripting languages (Python, JavaScript, TypeScript) versus systems programming languages (C, C++, C#, Java)?
RQ3: Scalability and Reliability How does CodeWiki’s hierarchical decomposition approach perform across repositories of varying sizes and complexity, and what is the reliability of our evaluation methodology?
4.2 Baselines
We compare CodeWiki against multiple repository-level documentation systems spanning both open-source and closed-source solutions:
Open-Source Baselines
OpenDeepWiki and deepwiki-open are open-source alternatives applying LLMs to entire repositories for documentation generation. These systems represent the current state of open-source repository-level documentation tools.
Closed-Source Baseline
DeepWiki (DeepWiki, 2025) is a commercial system has demonstrated effectiveness in industrial applications for automated repository-level documentation generation.
Positioning Relative to Function-Level Approaches
We note that function-level documentation systems such as RepoAgent (Luo et al., 2024) and DocAgent (Yang et al., 2025) generate N separate documents for N components, representing a fundamentally different documentation paradigm from CodeWiki’s hierarchically synthesized module-level outputs. Direct quantitative comparison would require forcing one approach into evaluation criteria it was not designed to satisfy, so we focus our comparison on systems targeting similar repository-level documentation goals.
4.3 Experimental Setup
Repository Selection
We curate a benchmark of 7 open-source repositories spanning diverse programming languages, project scales, and application domains. Our selection prioritizes: (1) language diversity for cross-language generalization, (2) varying codebase sizes ranging from 86K to 1.4M LOC, (3) actively maintained projects with high-quality official documentation for rubric construction, and (4) coverage of multiple domains such as web frameworks, automation tools, and machine learning platforms.
Table 1 presents the repository characteristics, including primary languages and codebase sizes. The benchmark covers JavaScript (svelte), TypeScript (puppeteer), C (wazuh), C++ (electron), C# (ml-agents), Java (logstash), and Python (OpenHands), ensuring comprehensive coverage across major programming paradigms.
Temporal Separation for Data Leakage Prevention
To ensure evaluation integrity, we use repository snapshots from August-September 2025, which postdate the knowledge cutoffs of all models used in our experiments (Claude Sonnet 4: March 2025). This temporal gap provides strong evidence against data leakage concerns. Detailed commit information is provided in Appendix E.
Model Configuration
Our experimental configuration employs state-of-the-art language models: Claude Sonnet 4 for documentation generation, multiple models (Claude Sonnet 4, Gemini 2.5 Pro, and Kimi K2 Instruct) for rubric generation to reduce single-model bias, and three distinct judge models (Gemini 2.5 Flash, GPT OSS 120B, and Kimi K2 Instruct) for independent assessments ensuring robust evaluation through multi-model consensus. Configuration details are in Appendix G.
Evaluation Protocol
For each repository, we execute the CodeWiki pipeline and generate comprehensive documentation including architectural overviews, component descriptions, and usage guides. The evaluation process follows our hierarchical rubric methodology: (1) generating repository-specific rubrics from official documentation, (2) performing agentic assessment of generated documentation, and (3) computing weighted scores with uncertainty quantification through standard deviation tracking.
4.4 Results and Analysis
4.4.1 RQ1: Documentation Quality and Coverage
Table 1 presents comprehensive evaluation results addressing the fundamental question of documentation quality. CodeWiki achieves an average documentation quality score of 68.79%, representing a 4.73 percentage point improvement over the closed-source DeepWiki baseline (64.06%) and substantial improvements over open-source alternatives OpenDeepWiki (47.13%) and deepwiki-open (50.05%). This demonstrates the effectiveness of our hierarchical synthesis approach compared to both simpler open-source methods and commercial solutions.
The coverage metrics reveal that CodeWiki consistently satisfies more leaf-level requirements than baselines in most cases. Notably, CodeWiki outperforms all baselines in 5 out of 7 repositories, with particularly strong improvements in TypeScript (83.00% vs. DeepWiki’s 64.46%), Python (82.45% vs. 73.04%), and JavaScript (71.96% vs. 68.51%). The open-source alternatives show notably degraded performance on more complex codebases, suggesting that simpler whole-repository prompting approaches do not scale effectively.
Answer: CodeWiki demonstrates superior overall documentation quality and coverage compared to both open-source and closed-source systems, achieving improvements of 21.66% over OpenDeepWiki, 18.74% over deepwiki-open, and 4.73% over DeepWiki.
4.4.2 RQ2: Language Generalization
Our analysis reveals distinct performance patterns across programming language categories. CodeWiki demonstrates strong performance on scripting languages, achieving average scores of 79.14% (Python, JavaScript, TypeScript) repositories - a 10.47% improvement over DeepWiki’s 68.67% average for the same languages.
For managed languages (C#, Java), CodeWiki achieves competitive performance with moderate but consistent improvements of 4.98% and 3.10% respectively, averaging 68.84% versus DeepWiki’s 64.80%. Systems programming languages (C, C++) present challenges where both frameworks exhibit similar patterns: DeepWiki achieves 56.39% while CodeWiki reaches 53.24%. This suggests that specialized parsing strategies may benefit both approaches for handling complex constructs like pointer and template metaprogramming.
Answer: CodeWiki demonstrates strong generalization with notable performance on scripting languages (79.14% average) and consistent improvements on managed languages, while systems programming languages present similar challenges for both frameworks.
4.4.3 RQ3: Scalability and Reliability
Repository size analysis demonstrates CodeWiki’s scalability across diverse project scales. Our framework processes repositories ranging from 86K LOC to 1.4M LOC, with performance consistency observed within language categories regardless of repository size. For high-level languages, CodeWiki maintains strong performance across different scales: Unity ML-Agents (86K LOC) achieves 79.78%, while the larger OpenHands (230K LOC) reaches 82.45%. This consistency demonstrates that our hierarchical approach successfully addresses context limitation challenges that typically constrain documentation generation at scale.
Critically, both CodeWiki and DeepWiki exhibit similar behavioral patterns across different repository sizes within the same programming language families. Systems programming languages present consistent challenges for both frameworks regardless of project scale, suggesting that the primary factor affecting documentation quality is language-specific parsing complexity rather than repository size. The dynamic delegation mechanism in CodeWiki automatically adapts to varying complexity levels, maintaining bounded processing requirements through hierarchical decomposition.
Reliability analysis through standard deviation measurements validates our multi-model assessment methodology. The average standard deviations of 3.84% (CodeWiki) versus 3.60% (DeepWiki) indicate comparable assessment reliability while achieving superior overall quality scores. A pilot human study (Appendix F) provides additional validation, showing that human preferences align with automated evaluation results (CodeWiki preferred in 7/9 assessments across 3 participants and 3 repositories).
Answer: CodeWiki’s hierarchical decomposition enables consistent scalability across repository sizes from 86K to 1.4M LOC, with performance variations primarily attributable to language-specific characteristics rather than scale limitations. Human evaluation validates alignment between our automated methodology and human judgment.
5 Conclusion
We present CodeWiki, an open-source framework for automated repository documentation using hierarchical decomposition and recursive processing. CodeWiki performs bottom-up hierarchical synthesis, generating parent modules from children while integrating visual artifacts. Across 7 multilanguage repositories, CodeWiki achieves 68.79% quality score, outperforming DeepWiki’s 64.06%. Performance variations correlate primarily with language characteristics rather than repository size, with scripting languages showing consistent improvements while systems languages face challenges for all frameworks.
Future work includes improved parsing for systems languages, version-aware documentation aligned with code evolution, and using documentation as semantic bridges for tasks like code migration. We release CodeWiki as open source to advance research in hierarchical documentation generation.
Limitations
While CodeWiki demonstrates strong performance across multiple languages and repository scales, several limitations warrant discussion.
Systems Programming Languages
Our framework exhibits reduced effectiveness on systems programming languages (C, C++) compared to high-level languages, achieving 53.24% versus 79.14% average scores. This gap stems from the inherent complexity of low-level constructs such as pointer operations, manual memory management, and template metaprogramming. For systems languages these constructs are the architecture, making this a limitation of our current analysis phase rather than a secondary concern. Developing specialized parsing modules that leverage language-specific characteristics represents an important direction for future work.
Rubric Generation
The automatically generated rubrics, while showing 73.65% semantic and 70.84% structural reliability across model families (Appendix D), have not undergone comprehensive human verification. Some generated rubrics may be difficult to interpret, and future work should incorporate systematic human validation to ensure rubric quality and clarity.
Evaluation Methodology
Our evaluation methodology relies on LLM-based assessment with multi-model consensus, which may introduce model-specific biases despite our mitigation strategies. A pilot human study (Appendix F) with 3 participants across 3 repositories showed alignment between human preferences and automated scores (CodeWiki preferred in 7/9 assessments), providing preliminary validation. However, broader human evaluation with more participants and systematic rubric-based assessment would further strengthen the benchmark’s validity. The average standard deviation of 3.84% indicates reasonable consistency, yet edge cases may benefit from expert validation.
Appendices
Appendix A Example Generated Documentation
Figure 3: Example repository-level documentation generated by CodeWiki for the All-Hands-AI–OpenHands repository. The documentation includes a comprehensive overview, architectural diagrams showing the modular event-driven architecture, and hierarchical navigation of core components including the agent system, implementations, and event processing modules.
Figure 3 shows an example of repository-level documentation generated by CodeWiki for the All-Hands-AI–OpenHands repository. The documentation provides a comprehensive overview including the repository’s purpose, end-to-end architecture with visual diagrams, and hierarchical navigation of core components. This demonstrates CodeWiki’s ability to synthesize both textual explanations and visual artifacts that capture architectural patterns and module relationships, enabling developers to quickly understand the system’s structure and design decisions.
Appendix B Background and Related Work
B.1 Code Summarization and Documentation
B.1.1 Existing Approaches
Advances in natural language processing (NLP) have enabled significant progress in the task of code summarization. Existing methods  (Khan and Uddin, 2023; Poudel et al., 2024; Makharev and Ivanov, 2025; Lomshakov et al., 2024; Ahmed and Devanbu, 2023; Choi et al., 2023; Luo et al., 2024; Yang et al., 2025) primarily focus on function-level documentation, and can broadly be divided into two categories based on the type of context they exploit (see Table 2 for a comprehensive comparison):
Local-Level Context
Early neural approaches established the foundation for automated documentation. CodeBERT (Feng et al., 2020) demonstrated effective bimodal training on code and natural language, while more recent LLM applications have shown substantial improvements. For instance, (Khan and Uddin, 2023) reported 20.6 BLEU scores across six languages using GPT-3, an 11.2% improvement over prior methods. DocuMint (Poudel et al., 2024) further explored the use of small language models for Python docstring generation with promising results. However, these approaches primarily rely on local code context and do not leverage repository-level information to improve documentation quality.
Codebase-Level Context
Recognizing the limitations of function-only context, recent work has explored broader codebase information to enrich documentation.  (Makharev and Ivanov, 2025) demonstrated that incorporating class- and repository-level context can substantially improve summary quality through few-shot learning and retrieval-augmented generation. ProConSuL (Lomshakov et al., 2024) incorporated project-level information via compiler analysis and call graphs, employing a two-phase training procedure to reduce hallucinations. RepoAgent (Luo et al., 2024) introduced comprehensive LLM-powered repository documentation, though it was limited in content diversity and architectural depth. DocAgent (Yang et al., 2025) advanced the field through multi-agent collaboration with topological processing, employing specialized agents that respect code dependencies to improve completeness and accuracy.
$\hookrightarrow$ Limitations and Our Motivation: Despite this progress, most approaches still target isolated snippets (e.g., functions or classes) and thus fail to capture the underlying architecture of the repository, as highlighted in Table 2. For a developer onboarding a completely new codebase, the primary challenge is not understanding what each individual function or class does, but rather grasping the high-level overview of the system, its dataflow and how different components interact. This underscores the need for an automatic methodology capable of generating documentation that spans from low-level details to high-level repository-wide perspectives, covering critical features and providing visual representations that support developer comprehension. Our approach, CodeWiki, addresses these limitations by generating true repository-level documentation while leveraging full codebase context and supporting multiple programming languages.
B.2 Repository-Level Documentation Systems
Several systems have addressed repository-level documentation generation with varying approaches and scopes:
Open-Source Alternatives
OpenDeepWiki and deepwiki-open are community-developed tools that apply LLMs to entire repositories. However, as shown in Table 1, these approaches exhibit degraded performance on larger codebases (averaging 47.13% and 50.05% respectively), suggesting that simpler whole-repository prompting does not scale effectively to complex projects.
RepoAgent (Luo et al., 2024) aggregates function and class-level documentation into repository structures using dependency analysis. While effective for component cataloging, it does not perform hierarchical synthesis where parent modules summarize and contextualize their children.
codebase2tutorial [^3] https://code2tutorial.com/ generates tutorial-style documentation using whole-repository prompting, focusing on educational walkthroughs rather than comprehensive architectural documentation.
LARCH (Koreeda et al., 2023) targets README generation specifically, addressing a narrower scope than full repository documentation.
Agent-as-a-Judge (Zhuge et al., 2024) includes an OpenWiki feature that leverages existing READMEs and applies LLM assessment, representing a different methodology focused on evaluation rather than generation.
CodeWiki’s Distinguishing Characteristics
CodeWiki differs from these approaches through: (1) true hierarchical synthesis where parent modules are generated by summarizing and contextualizing child documentation, (2) integrated visual artifacts (architecture diagrams, data flow visualizations) generated alongside textual content, (3) dynamic delegation enabling scalable processing of arbitrarily large repositories, and (4) cross-module reference management maintaining documentation coherence across module boundaries.
B.2.1 Existing Evaluation Methods
Most existing works evaluate the quality of generated documentation using textual similarity metrics such as BLEU or ROUGE. These metrics measure the overlap between generated text and reference ground-truth documentation in terms of lexical or token-level similarity.
$\hookrightarrow$ Limitations and Our Motivation: While these methods provide a lightweight quantitative measure, they often fail to capture deeper semantic aspects of the documentation, such as clarity, correctness, and completeness in explaining the underlying code behavior. Recent studies (Roy et al., 2021; Evtikhiev et al., 2023) have shown that traditional metrics like BLEU and ROUGE do not consistently correlate with human judgment of documentation quality. This gap highlights the need for an assessment system that can fairly and accurately evaluate documentation, ensuring that the generated content truly supports developers in comprehending the system.
Table 2: Comparison of different code summarization approaches. Repo-Level Doc. indicates whether the approach generates repository-level documentation output. Codebase Context specifies whether the codebase information is used as input to generate the documentation.
Approach Repo-Level Doc. Codebase Context Multilingual DocuMint ASAP ProConSul DocAgent RepoAgent CodeWiki
B.3 Rubric-based evaluation
A repository can be described through multiple representations, each using different structures and presentation styles while sharing the same underlying architecture. This diversity makes it difficult to evaluate open-ended outputs automatically, as no single metric can reliably evaluate which version is better. Although human evaluation provides trustworthy results, it is resource-intensive and slow. To overcome these limitations and facilitate rapid assessment, many benchmarks (Starace et al., 2025; Lin et al., 2025; Sirdeshmukh et al., 2025; Arora et al., 2025) often employ large language models (LLMs) as automated evaluators, using task-specific rubrics to judge the quality of generated outputs. In our CodeWikiBench benchmark, we similarly construct a set of hierarchical rubrics based on the official documentation, providing a structured and fine-grained framework to evaluate the quality and completeness of generated repository-level documentation.
Appendix C Algorithm Details
Input: Module tree $T$, dependency graph $G$
Output: Complete repository documentation
1 Initialize documentation workspace;
2 for each leaf module $m$ in $T$ do
3   agent $\leftarrow$ CreateAgent($m$, $G$, $T$);
4 while $m$ not fully processed do
5      agent.GenerateDoc($m$.name);
6    if delegation requested then
7         sub $\leftarrow$ agent.GetDelegationSpecs();
8         UpdateModuleTree($T$, $m$, sub);
9       foreach new leaf $m^{\prime}$ in sub do
10            Recursively process $m^{\prime}$;
11
12         end foreach
13        ReviseDoc($m$.name);
14
15      end if
16
17   end while
18
19 end for
20ProcessParentModules($T$);
21 return Repository documentation;
Algorithm 1 Recursive Module Documentation
Appendix D Rubric Reliability Assessment
To ensure the quality and trustworthiness of automatically generated evaluation rubrics, we developed a comprehensive reliability assessment framework that quantifies the consistency and structural integrity of rubrics produced by different language models. This assessment methodology provides empirical measures of rubric reliability across multiple dimensions.
Given a set of $n$ rubrics $\{R_{1},R_{2},\ldots,R_{n}\}$ produced by different models for the same documentation corpus, we compute pairwise consistency scores across both semantic and structural dimensions.
D.1 Semantic Consistency Analysis
We evaluate semantic consistency by extracting all requirement texts from each rubric and compute semantic similarity using distributed representations. Let $T_{i}=\{t_{1}^{(i)},t_{2}^{(i)},\ldots,t_{m_{i}}^{(i)}\}$ denote the set of requirement texts extracted from rubric $R_{i}$, where $m_{i}$ represents the total number of requirements in $R_{i}$.
For each requirement text $t_{k}^{(i)}\in T_{i}$, we identify its best semantic match in $T_{j}$ using cosine similarity between embeddings:
$\text{sim}_{\text{best}}(t_{k}^{(i)},T_{j})=\max_{t_{\ell}^{(j)}\in T_{j}}\cos(\mathbf{e}(t_{k}^{(i)}),\mathbf{e}(t_{\ell}^{(j)}))$ (1)
where $\mathbf{e}(\cdot)$ represents the embedding function and $\cos(\cdot,\cdot)$ denotes cosine similarity. The bidirectional semantic similarity between rubrics is then:
$\text{Sim}_{\text{semantic}}(R_{i},R_{j})$ $=\frac{1}{m_{i}+m_{j}}\Bigg(\sum_{k=1}^{m_{i}}\text{sim}_{\text{best}}(t_{k}^{(i)},T_{j})$
$\quad+\sum_{\ell=1}^{m_{j}}\text{sim}_{\text{best}}(t_{\ell}^{(j)},T_{i})\Bigg)$ (2)
This formulation ensures that both the coverage of requirements from $R_{i}$ in $R_{j}$ and vice versa are considered, providing a balanced measure of semantic overlap.
D.2 Structural Consistency Analysis
Structural consistency evaluates the architectural similarity between rubrics, focusing on hierarchical organization rather than content semantics. We define a comprehensive set of structural features for each rubric:
  • Maximum depth ($d$): The deepest level in the rubric hierarchy
  • Total items ($N$): The total number of rubric items across all levels
  • Weight distribution ($W$): The frequency distribution of importance weights
The structural similarity between rubrics $R_{i}$ and $R_{j}$ combines multiple normalized metrics:
$\text{Sim}_{\text{depth}}(R_{i},R_{j})=1-\frac{|d_{i}-d_{j}|}{\max(d_{i},d_{j},1)}$ (3)
$\text{Sim}_{\text{items}}(R_{i},R_{j})=1-\frac{|N_{i}-N_{j}|}{\max(N_{i},N_{j},1)}$ (4)
For weight distribution similarity, we compute the overlap between normalized probability distributions:
$\text{Sim}_{\text{weights}}(R_{i},R_{j})=\sum_{w\in W_{i}\cup W_{j}}\min(P_{i}(w),P_{j}(w))$ (5)
where $P_{i}(w)$ represents the normalized frequency of weight $w$ in rubric $R_{i}$.
The overall structural similarity is computed as:
$\text{Sim}_{\text{structural}}(R_{i},R_{j})$ $=\frac{1}{3}\Big(\text{Sim}_{\text{depth}}(R_{i},R_{j})$
$\quad+\text{Sim}_{\text{items}}(R_{i},R_{j})$ (6)
$\quad+\text{Sim}_{\text{weights}}(R_{i},R_{j})\Big)$ (7)
D.3 Overall Reliability Score
The overall reliability score aggregates consistency measures across all model pairs. For $n$ models, we compute $\binom{n}{2}$ pairwise comparisons and derive summary statistics:
$\text{Score}^{semantic}_{\text{reliability}}=\frac{1}{\binom{n}{2}}\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}\text{Sim}_{\text{semantic}}(R_{i},R_{j})$ (8)
$\text{Score}^{structural}_{\text{reliability}}=\frac{1}{\binom{n}{2}}\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}\text{Sim}_{\text{structural}}(R_{i},R_{j})$ (9)
Table 3 presents the reliability assessment results across our benchmark repositories, demonstrating consistent rubric generation with average semantic reliability of 73.65% and structural reliability of 70.84%.
Table 3: Rubric Reliability Assessment: Semantic and Structural Consistency Scores
Repository Semantic (%) Structural (%) All-Hands-AI–OpenHands 73.64 68.82 sveltejs–svelte 75.56 68.88 puppeteer–puppeteer 72.97 76.36 Unity-Technologies–ml-agents 75.95 71.22 elastic–logstash 73.13 68.22 wazuh–wazuh 72.20 67.62 electron–electron 72.12 74.79 Average 73.65 70.84
Appendix E Temporal Separation Details
Table 4 presents the specific commit information for each benchmark repository, demonstrating temporal separation from model training data.
Table 4: Repository commit information demonstrating temporal separation from model training data.
Repository Language Commit ID Commit Date OpenHands Python 30604c4 Sep 13, 2025 svelte JavaScript be645b4 Aug 31, 2025 puppeteer TypeScript c1105f1 Aug 29, 2025 ml-agents C# 4cf2f49 Aug 15, 2025 logstash Java 895cfa5 Aug 29, 2025 wazuh C 44b7cd3 Aug 30, 2025 electron C++ 828fd59 Aug 30, 2025
Appendix F Human Evaluation Pilot Study
To provide additional validation of our automated evaluation methodology beyond multi-model consensus, we conducted a pilot human study comparing CodeWiki and DeepWiki documentation outputs.
F.1 Study Design
Participants
We recruited three participants with diverse professional backgrounds relevant to documentation consumption:
  • Business Analyst: Represents stakeholders seeking high-level architectural understanding
  • AI Researcher: Represents technical users evaluating documentation for research purposes
  • Software Engineer: Represents developers using documentation for implementation guidance
None of the participants were prior contributors to the evaluated repositories, simulating the common scenario where developers approach unfamiliar codebases.
Evaluation Protocol
For each of three repositories (OpenHands, svelte, puppeteer), participants:
  1. 1.
    Received access to the official codebase and its official documentation
  2. 2.
    Reviewed anonymized documentation outputs from both CodeWiki and DeepWiki (presentation order randomized)
  3. 3.
    Assessed which documentation better reflected: (a) correctness relative to official documentation, and (b) coverage of actual repository architecture
  4. 4.
    Provided a binary preference judgment
The evaluation was conducted blind, with system identities hidden from participants.
F.2 Results
Table 5: Pilot human evaluation results showing participant preferences between CodeWiki (CW) and DeepWiki (DW) across three repositories.
Repository BA AI SE CW Pref.
OpenHands CW CW CW 3/3
svelte CW CW DW 2/3
puppeteer CW DW CW 2/3
Total 3/3 2/3 2/3 7/9 (77.8%)
Table 5 presents the evaluation results. CodeWiki was preferred in 7 out of 9 assessments (77.8%). Key observations:
Alignment with Automated Scores
Human preferences correlate with the magnitude of automated evaluation improvements:
  • OpenHands (+9.41% automated improvement): Unanimous CodeWiki preference (3/3)
  • puppeteer (+18.54% automated improvement): Strong CodeWiki preference (2/3)
  • svelte (+3.45% automated improvement): Closer preferences (2/3 for CodeWiki)
This alignment between human judgment and automated evaluation supports the validity of our LLM-based assessment methodology.
F.3 Limitations
We acknowledge several limitations of this pilot study:
  • Scale: Three participants and three repositories provide preliminary validation but not statistical significance
  • Simplified Protocol: Due to budget constraints, we used preference judgment rather than full rubric-based evaluation
  • Repository Selection: The evaluated repositories may not represent the full diversity of our benchmark
Comprehensive human evaluation with larger participant pools, more repositories, and detailed rubric-based assessment represents important future work to further validate the benchmark methodology.
Appendix G Implementation Details
G.1 Model Configurations
Documentation Generation Agent
  • Model: Claude Sonnet 4 (claude-sonnet-4-20250514)
  • Temperature: 0.0
  • Max tokens: 16384
  • Context window: 200K tokens
Rubric Generation Agents
  • Model 1: Claude Sonnet 4 (Temperature: 0.0)
  • Model 2: Gemini 2.5 Pro (Temperature: 0.0)
  • Model 3: Kimi K2 Instruct (Temperature: 0.0)
  • Max tokens per generation: 32768
Judge Agents
  • Judge 1: Gemini 2.5 Flash (Temperature: 0.0)
  • Judge 2: GPT OSS 120B (Temperature: 0.0)
  • Judge 3: Kimi K2 Instruct (Temperature: 0.0)
  • Binary scoring: 0 (not satisfied) or 1 (satisfied)
G.2 Processing Parameters
  • Module decomposition threshold: 32768 tokens per leaf module
  • Maximum delegation depth: 3 levels